# Localization and interpolation of parabolic $L^{p}$ Neumann problems

### v1

@@ 2.2. Reinforced weak and Energy solutions
and therefore defines a bounded linear functional on $\dot{\mathsf{E}}_{0}$. This again implies the existence of a solution for the Poisson Dirichlet problem with righthand side containing the term $h$. We do not consider the Neumann problem for the term $h$.
@@ 2.3. The $L^{p}$ Poisson problems
$\|N(\nabla u)\|_{L^{p}(\partial\Omega)}\leq C\left(\|\widetilde{\mathcal{C}}(\delta h)\|_{L^{p}(\partial\Omega)}+\|\widetilde{\mathcal{C}}(F)\|_{L^{p}(\partial\Omega)}\right).$
@@
Let $p\in(1,\infty)$. We say that the weak $L^{p}$ Poisson-Neumann-Regularity $\rm{(}\rm{wPNR}_{p}\rm{)}$ is solvable for $L$ if for any ${F}\in L^{\infty}_{c}(\Omega,\mathbb{R}^{n})$, the solution the solution $u$ to $Lu=-\operatorname{div}F$ with 0 Neumann boundary data satisfies
@@ 3. Basic estimates
Let $Q=Q_{r}(X,t)$ be a parabolic cube that either satisfies $4Q=Q_{4r}(X,t)\subset\Omega$ or $(X,t)\in\partial\Omega$ (i.e. centered at $\partial\Omega$). Let $u$ be a weak solution of $Lu=0$ in $Q_{4r}(X,t)\cap\Omega$, and assume that $u$ has zero Neumann data on $4Q\cap\partial\Omega$ if $(X,t)\in\partial\Omega$. Then there exists a constant $C=C(\lambda,\Lambda,n)$ such that for any constant $c$,
$\iint_{Q_{r}(X,t)\cap\Omega}|\nabla u|^{2}dYds\leq\frac{C}{r^{2}}\iint_{Q_{2r}(X,t)\cap\Omega}|u(Y,s)-c|^{2}dYds.$
@@
The proof of the interior Caccioppoli is classical. For the case when $Q$ is centered at $\partial\Omega$, we assume that $\mathcal{O}\cap 4Q=\left\{(x,x_{n})\in\mathbb{R}^{n}:x_{n}>\varphi(x)\right\}\cap 4Q$ for some Lipschitz function $\varphi$, and consider the even reflection of $u$ defined as
@@
Since $u$ has zero Neumann data on $4Q\cap\partial\Omega$, $\tilde{u}$ is a weak solution to $\tilde{L}\tilde{u}=-\partial_{t}\tilde{u}+\operatorname{div}(\tilde{A}\nabla\tilde{u})=0$ in $4Q$, where $\tilde{A}=A$ on $\Omega$ and the coefficient matrix $\tilde{A}$ on the set $x_{n}<\varphi(x)$ depends only on $A(x,2\varphi(x)-x_{n},t)$ and $\nabla\varphi$ making the resulting matrix bounded and uniformly elliptic on the set $4Q\setminus\Omega$. Applying the interior Caccioppoli’s inequality to $\tilde{u}$ gives
$\iint_{Q}|\nabla\tilde{u}|^{2}dYds\leq\frac{C}{r^{2}}\iint_{2Q}|\tilde{u}(Y,s)-c|^{2}dYds,$
@@
$\iint_{Q\cap\Omega}|\nabla u|^{2}dYds\lesssim\iint_{Q}|\nabla\tilde{u}|^{2}dYds,\text{ and }\iint_{2Q}|\tilde{u}(Y,s)-c|^{2}dYds\lesssim\iint_{2Q\cap\Omega}|u(Y,s)-c|^{2}dYds.$
@@
Let $Q$ and $u$ be as in Lemma 3.1. There exists constant $C=C(\lambda,\Lambda,n)$ such that
(3.4) $\sup_{Q_{r/2}(X,t)\cap\Omega}|u|\leq\frac{C}{r^{n+2}}\iint_{Q_{2r}(X,t)\cap\Omega}|u(Y,s)|dYds.$
@@
When $4Q\subset\Omega$, i.e. in the interior case, the result is classical. See for example [1, Theorem B], which gives (3.4) with an $L^{2}$ version on the right-hand side. Going from $L^{2}$ to $L^{1}$ is also classical and is based on a dilation argument. When $4Q$ is centered at the boundary, we use the reflection (3.2) to reduce matters to the interior case. We skip the details. ∎
@@
Let $Q=Q_{r}(X,t)$ be a parabolic cube that either satisfies $4Q=Q_{4r}(X,t)\subset\Omega$ or $(X,t)\in\partial\Omega$ (i.e. centered at $\partial\Omega$). Let $u$ be a weak solution of $Lu=-\mathrm{div}F$ in $Q_{4r}(X,t)\cap\Omega$, and assume that $u$ has zero Neumann data on $4Q\cap\partial\Omega$ if $(X,t)\in\partial\Omega$. Then there is some constant $C$ such that
$\iint_{Q\cap\Omega}|u-c_{u}|dXdt\leq Cr\iint_{2Q\cap\Omega}\left(|\nabla u|+|F|\right)dXdt,$
where $c_{u}=\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{Q\cap\Omega}u\,dXdt$.
@@
We only prove the case when $Q$ is centered at $\partial\Omega$, as the interior case is similar. We can assume without loss of generality that $Q$ is centered at $(0,0)\in\partial\Omega$. We write $Q=Q^{\prime}\times[-r^{2},r^{2}]$, where $Q^{\prime}=\left\{X\in\mathbb{R}^{n}:|x_{i}|<r\right\}$ is the cube in $\mathbb{R}^{n}$ centered at 0 with sidelength $r$. For $t\in\mathbb{R}$, we set $\langle u\rangle(t):=\fint_{Q^{\prime}\cap\mathcal{O}}u(Y,t)dY$. By the triangle inequality and the Poincaré inequality in the spatial variables,
(3.6) $\iint_{Q\cap\Omega}|u(X,t)-c_{u}|dXdt\leq\iint_{Q\cap\Omega}|u(X,t)-\langle u\rangle(t)|dXdt+\iint_{Q\cap\Omega}|\langle u\rangle(t)-c_{u}|dXdt\\ \leq Cr\int_{-r^{2}}^{r^{2}}\int_{Q^{\prime}\cap\mathcal{O}}\left|\nabla u(X,t)\right|dXdt+|Q^{\prime}\cap\mathcal{O}|\int_{-r^{2}}^{r^{2}}|\langle u\rangle(t)-c_{u}|dt.$
@@
$\langle u\rangle_{\eta}(t):=\int_{\mathcal{O}}u(Y,t)\eta(Y)dY\quad\text{for }t\in\mathbb{R},$
and set $c_{\eta}:=\int_{\mathcal{O}}\eta\,dX$. For $t,t^{\prime}\in[-r^{2},r^{2}]$, we have
@@
since $u$ is a weak solution to $Lu=-\operatorname{div}F$ with zero Neumann data on $\partial\Omega\cap 4Q$. Using the bound on $\nabla\eta$, we obtain that
(3.7) $|\langle u\rangle_{\eta}(t)-\langle u\rangle_{\eta}(t^{\prime})|\leq\frac{C}{r}\int_{\mathcal{O}\cap 2Q^{\prime}}\int_{-r^{2}}^{r^{2}}\left(|\nabla u|+|F|\right)dtdX\quad\text{for }t,t^{\prime}\in[-r^{2},r^{2}].$
@@
$\int_{-r^{2}}^{r^{2}}|\langle u\rangle(t)-c_{u}|dt=\int_{-r^{2}}^{r^{2}}\left|\fint_{-r^{2}}^{r^{2}}\left(\langle u\rangle(t)-\langle u\rangle(t^{\prime})\right)dt^{\prime}\right|dt\leq Cr\fint_{\mathcal{O}\cap 2Q^{\prime}}\int_{-r^{2}}^{r^{2}}\left(|\nabla u|+|F|\right)dtdX.$
@@
$\iint_{Q\cap\Omega}|u(X,t)-c_{u}|dXdt\leq Cr\int_{\mathcal{O}\cap 2Q^{\prime}}\int_{-r^{2}}^{r^{2}}\left(|\nabla u|+|F|\right)dtdX.$
In the following lemma we show relations between the Dirichlet and Poisson Dirichlet problems. We remark that although it can be shown that the solvability of $\rm{(}\rm{D}_{p^{\prime}}\rm{)}$$^{L^{*}}$, $\rm{(}\rm{PD}_{p^{\prime}}\rm{)}$$^{L^{*}}$ and $\rm{(}\rm{PR}_{p}\rm{)}$L are actually equivalent, we will only derive the partial implications that are sufficient for our purpose.
@@
by the assumption that $\rm{(}\rm{PD}_{p^{\prime}}\rm{)}$$^{L^{p^{\prime}}}$ is solvable. Again by (2.13), we have
@@
(4.3) $\|\widetilde{N}(\nabla u)\|_{L^{p}(\partial\Omega)}\leq\|\widetilde{N}(\nabla u_{D})\|_{L^{p}(\partial\Omega)}+\|\widetilde{N}(\nabla w)\|_{L^{p}(\partial\Omega)}\\ \leq|\widetilde{N}(\nabla u_{D})\|_{L^{p}(\partial\Omega)}+\|g\|_{L^{p}(\partial\Omega)}\leq 2\|\widetilde{N}(\nabla u_{D})\|_{L^{p}(\partial\Omega)}\lesssim\|\widetilde{\mathcal{C}}(F)\|_{L^{p}(\partial\Omega)}.$
@@
where we have used that $\|F\|_{L^{2}(\Omega)}<\infty$ since $F\in L_{c}^{\infty}(\Omega)$. Hence, there exists a unique $g\in\dot{H}{\vphantom{H}}^{-1/4}_{\partial_{t}-\Delta_{x}}(\partial\Omega)$ such that $\Lambda(\varphi)=\langle g,\varphi\rangle$ for all $\varphi\in\dot{H}^{1/4}_{\partial_{t}-\Delta_{x}}(\partial\Omega)$. As Lipschitz functions $\varphi:\mathbb{R}^{n+1}\times\mathbb{R}$ with bounded support restricted to $\partial\Omega$ are dense in $\dot{H}^{1/4}_{\partial_{t}-\Delta_{x}}(\partial\Omega)$ it follows that the identity $\Lambda(\varphi)=\langle g,\varphi\rangle$ holds for all such $\varphi$ (and we can take $\Phi$ to be their Lispchitz extension in (4.5)) and still uniquely determined $g$ in the space $\dot{H}{\vphantom{H}}^{-1/4}_{\partial_{t}-\Delta_{x}}(\partial\Omega)$.
@@
To show the weak convergence of $g_{r}$ to $g$ in $L^{p}(\partial\Omega)$, it suffices to show that for any Lipschitz function $\varphi:\mathbb{R}^{n+1}\times\mathbb{R}$ with bounded support, we have that
@@
(4.14) $I\leq\iint_{\Omega}u\,Lv\,dX=\iint_{\Omega}u\left(\partial_{t}v+\operatorname{div}A\nabla v\right)dXdt=\iint_{\Omega}\left(\partial_{t}u\,v-A^{T}\nabla u\cdot\nabla v\right)\,dXdt\\ =\iint_{\Omega}F\cdot\nabla v\,dXdt$
because $u$ is a weak solution to $L^{*}u=-\mathrm{div}F$ with zero Neumann data. By the Carleson inequality (2.13), $\rm{(}\rm{wPNR}_{p}\rm{)}$ holds for $v$, and (4.13), we have that
@@
Let $p\in(1,\infty)$. Suppose that $\rm{(}\rm{wPN}_{p^{\prime}}\rm{)}$$^{L^{*}}$ is solvable in $\Omega=\mathcal{O}\times\mathbb{R}$. Then for any parabolic cube $Q=Q_{r}(x,t)$ centered at some $(x,t)\in\partial\Omega$, and any local solution $u\in W^{1,2}(2Q\cap\Omega)$ to $Lu=0$ in $2Q\cap\Omega$ with zero Neumann data on $2Q\cap\partial\Omega$, we have
(4.16) $\|\widetilde{N}(|\nabla u|{\mathds{1}}_{Q})\|_{L^{p}(\partial\Omega)}\leq Cr^{(n+1)/p}\left(\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{2Q\cap\Omega}|\nabla u|^{2}dXdt\right)^{\frac{1}{2}}.$

### v2

@@ 2.2. Reinforced weak and Energy solutions
and therefore defines a bounded linear functional on $\dot{\mathsf{E}}_{0}$. This again implies the existence of a solution for the Poisson Dirichlet problem with right-hand side containing the term $h$. We do not consider the Neumann problem for the term $h$.
More on weak solutions. Let $u$ be a weak solution to $Lu=0$ in $\Omega$, that is, $u\in L_{\operatorname{loc}}^{2}(\mathbb{R};\mathcal{W}^{1,2}_{\operatorname{loc}}(\mathcal{O}))$, $\partial_{t}u\in L_{\operatorname{loc}}^{2}(\mathbb{R};\mathcal{W}^{-1,2}_{\operatorname{loc}}(\mathcal{O}))$, and (2.18) holds for all $\phi\in C_{c}^{\infty}(\Omega)$. We can rewrite (2.18) as
(2.24) $\iint_{\Omega}A\nabla u\cdot{\nabla\phi}\,\mathrm{d}X\,\mathrm{d}t+\int_{\mathbb{R}}\langle\partial_{t}u,\phi\rangle\,\mathrm{d}t=0,$
where the pairing $\langle\partial_{t}u,\phi\rangle$ is understood as pairing of $\mathcal{W}^{-1,2}_{\operatorname{loc}}(\mathcal{O})$ with $\mathcal{W}^{1,2}_{0}(\mathcal{O})$ function subsequently integrated in time. We refer to solutions $u$ such that $u\in L_{\operatorname{loc}}^{2}((t_{0},t_{1});\mathcal{W}^{1,2}_{\operatorname{loc}}(\mathcal{O}))$ and $\partial_{t}u\in L_{\operatorname{loc}}^{2}((t_{0},t_{1});\mathcal{W}^{-1,2}_{\operatorname{loc}}(\mathcal{O}))$ when the support of $\phi$ is limited to the set $\mathcal{O}\times[t_{0},t_{1}]$.
Let $\theta_{\varepsilon}:\mathbb{R}\to\mathbb{R}$ be a family of standard mollifiers, that is, $\theta_{\varepsilon}(\tau)=\frac{1}{\varepsilon}\theta(\tau/\varepsilon)$ for $\tau\in\mathbb{R}$, $\varepsilon>0$, where $\theta\in C_{c}^{\infty}(-1,1)$, $\int_{\mathbb{R}}\theta=1$, and $0\leq\theta\leq 1$. Let $\psi\in C_{c}^{\infty}(\mathcal{O})$ be a fixed function. For a fixed $t\in(t_{0},t_{1})$, we consider the smooth function $\phi_{\varepsilon}(X,\tau):=\psi(X)\theta_{\varepsilon}(t-\tau)$. Then for a weak solution $u$ on $\mathcal{O}\times(t_{0},t_{1})$, thanks to (2.24) we have that
(2.25) $\iint_{\Omega}A(X,\tau)\nabla u(X,\tau)\cdot{\nabla\psi(X)}\,\theta_{\varepsilon}(t-\tau)\,\mathrm{d}X\,\mathrm{d}\tau+\int_{\mathbb{R}}\langle\partial_{\tau}u(\cdot,\tau),\psi\rangle\,\theta_{\varepsilon}(t-\tau)\,\mathrm{d}\tau=0.$
As $\varepsilon\to 0_{+}$, the first term of (2.25) converges to
$\int_{\mathcal{O}}A(X,t)\nabla u(X,t)\cdot{\nabla\psi(X)}\,\mathrm{d}X\qquad\text{for a.e. }t\in(t_{0},t_{1}).$
For the second term, since $\partial_{t}u\in L_{\operatorname{loc}}^{2}((t_{0},t_{1});\mathcal{W}^{-1,2}_{\operatorname{loc}}(\mathcal{O}))$, the function $\tau\mapsto\langle\partial_{\tau}u(\cdot,\tau),\psi\rangle$ is in $L_{\operatorname{loc}}^{2}(t_{0},t_{1})$, and therefore, for a.e. $t\in(t_{0},t_{1})$,
$\int_{\mathbb{R}}\langle\partial_{\tau}u(\cdot,\tau),\psi\rangle\,\theta_{\varepsilon}(t-\tau)\,\mathrm{d}\tau\to\langle\partial_{\tau}u(\cdot,t),\psi\rangle\qquad\text{as }\varepsilon\to 0_{+}.$
Hence for a.e. $t\in(t_{0},t_{1})$, we have
(2.26) $\langle\partial_{\tau}u(\cdot,t),\psi\rangle+\int_{\mathcal{O}}A(X,t)\nabla u(X,t)\cdot{\nabla\psi(X)}\,\mathrm{d}X=0\qquad\text{for any }\psi\in C_{c}^{\infty}(\mathcal{O}),$
and equivalently (by a density argument), for any $\psi\in\mathcal{W}_{0}^{1,2}(\mathcal{O})$. In particular, it implies that if $u$ is a weak solution to $Lu=0$ in $\mathcal{O}\times(t_{0},t_{1})$, then
(2.27) $\int_{t_{0}}^{t_{1}}\langle\partial_{\tau}u(\cdot,t),\phi(\cdot,t)\rangle\,dt+\int_{t_{0}}^{t_{1}}\int_{\mathcal{O}}A\nabla u\cdot{\nabla\phi}\,\mathrm{d}Xdt=0\quad\text{for any }\phi\in L_{\operatorname{loc}}^{2}((t_{0},t_{1});\mathcal{W}_{0}^{1,2}(\mathcal{O})),$
since we can take $\psi=\phi(\cdot,t)$ in (2.26) and integrate in $t$ over $(t_{0},t_{1})$. The observations (2.26) and (2.27) will be very useful since they allow us to take sharp cut-offs in time.
We remark that similar observations also apply to weak solutions (and therefore, energy solutions) to the Poisson problem $Lu=-\mathrm{div}F$ with zero Dirichlet and Neumann boundary data; we skip the details.
@@ 2.3. The $L^{p}$ Poisson problems
$\|\widetilde{N}(\nabla u)\|_{L^{p}(\partial\Omega)}\leq C\left(\|\widetilde{\mathcal{C}}(\delta h)\|_{L^{p}(\partial\Omega)}+\|\widetilde{\mathcal{C}}(F)\|_{L^{p}(\partial\Omega)}\right).$
@@
Let $p\in(1,\infty)$. We say that the weak $L^{p}$ Poisson-Neumann-Regularity $\rm{(}\rm{wPNR}_{p}\rm{)}$ is solvable for $L$ if for any ${F}\in L^{\infty}_{c}(\Omega,\mathbb{R}^{n})$, the solution $u$ to $Lu=-\operatorname{div}F$ with 0 Neumann boundary data satisfies
@@ 3. Basic estimates
Let $Q_{0}\subset\mathbb{R}^{n+1}$ be a cube satisfying $Q_{0}\cap\Omega\neq\emptyset$. Let $u$ be a weak solution of $Lu=0$ in $Q_{0}\cap\Omega$. If $Q_{0}\cap\partial\Omega\neq\emptyset$, then we assume in addition that $u$ has zero Neumann data on ${Q_{0}}\cap\partial\Omega$.
Let $(\bar{X},\bar{t})\in Q_{0}\cap\overline{\Omega}$ and let $J_{r}:=J_{r}(\bar{X},\bar{t})$ be a backward in time parabolic cube that satisfies $J_{4r}\subset Q_{0}$. Then there exists a constant $C=C(\lambda,\Lambda,n)$ such that for any constant $c$,
$\iint_{J_{r}\cap\Omega}|\nabla u|^{2}dYds\leq\frac{C}{r^{2}}\iint_{J_{2r}\cap\Omega}|u(Y,s)-c|^{2}dYds.$
@@
When $Q_{0}\subset\Omega$, it is the standard interior Caccioppoli inequality whose proof is classical. When $Q_{0}\cap\partial\Omega\neq\emptyset$, we assume without loss of generality that it is centered at $\partial\Omega$. Assume that $\mathcal{O}\cap Q_{0}=\left\{(x,x_{n})\in\mathbb{R}^{n}:x_{n}>\varphi(x)\right\}\cap Q_{0}$ for some Lipschitz function $\varphi$, and consider the even reflection of $u$ defined as
@@
Since $u$ has zero Neumann data on $Q_{0}\cap\partial\Omega$, $\tilde{u}$ is a weak solution to $\tilde{L}\tilde{u}=-\partial_{t}\tilde{u}+\operatorname{div}(\tilde{A}\nabla\tilde{u})=0$ in $Q_{0}$, where $\tilde{A}=A$ on $\Omega$ and the coefficient matrix $\tilde{A}$ on the set $x_{n}<\varphi(x)$ depends only on $A(x,2\varphi(x)-x_{n},t)$ and $\nabla\varphi$ making the resulting matrix bounded and uniformly elliptic on the set $Q_{0}\setminus\Omega$. Applying the interior Caccioppoli’s inequality to $\tilde{u}$ gives
$\iint_{J_{r}}|\nabla\tilde{u}|^{2}dYds\leq\frac{C}{r^{2}}\iint_{J_{2r}}|\tilde{u}(Y,s)-c|^{2}dYds,$
@@
$\iint_{J_{r}\cap\Omega}|\nabla u|^{2}dYds\lesssim\iint_{J_{r}}|\nabla\tilde{u}|^{2}dYds,\text{ and }\iint_{J_{2r}}|\tilde{u}(Y,s)-c|^{2}dYds\lesssim\iint_{J_{2r}\cap\Omega}|u(Y,s)-c|^{2}dYds.$
@@
Let $Q_{0}$, $J_{r}$, and $u$ be as in Lemma 3.1. There exists a constant $C=C(\lambda,\Lambda,n)$ such that
(3.4) $\sup_{J_{r/2}\cap\Omega}|u|\leq\frac{C}{r^{n+2}}\iint_{J_{2r}\cap\Omega}|u(Y,s)|dYds.$
@@
When $Q_{0}\subset\Omega$, i.e., in the interior case, the result is classical. See for example [1, Theorem B], which gives (3.4) with an $L^{2}$ version on the right-hand side. Going from $L^{2}$ to $L^{1}$ is also classical and is based on a dilation argument. When $Q_{0}\cap\partial\Omega\neq\emptyset$, we use the reflection (3.2) to reduce matters to the interior case. We skip the details. ∎
@@
Let $Q_{0}=Q_{0}^{\prime}\times(T_{0},T_{1})$ be a cube satisfying $Q_{0}\cap\Omega\neq\emptyset$. Let $u$ be a weak solution of $Lu=-\mathrm{div}F$ in $Q_{0}\cap\Omega$, and assume in addition that $u$ has zero Neumann data on $Q_{0}\cap\partial\Omega$ if $Q_{0}\cap\partial\Omega\neq\emptyset$. Let $(X,t)\in Q_{0}\cap\overline{\Omega}$ and let $J_{r}:=J_{r}(X,t)$ be a backward in time parabolic cube that satisfies $J_{4r}\subset Q_{0}$.
Then there is some constant $C$ such that
$\iint_{J_{r}\cap\Omega}|u-c_{u}|dXdt\leq Cr\iint_{J_{2r}\cap\Omega}\left(|\nabla u|+|F|\right)dXdt,$
where $c_{u}=\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{J_{r}\cap\Omega}u\,dXdt$.
@@
We only prove the case when $Q_{0}\cap\partial\Omega\neq\emptyset$, as the interior case is similar. We can assume without loss of generality that $(0,0)\in Q_{0}\cap\partial\Omega$ and $J_{r}=Q^{\prime}\times(-r^{2},0)$, where $Q^{\prime}=\left\{X\in\mathbb{R}^{n}:|x_{i}|<r\right\}$ is the cube in $\mathbb{R}^{n}$ centered at 0 with sidelength $r$. For $t\in(T_{0},T_{1})$, we set $\langle u\rangle(t):=\fint_{Q^{\prime}\cap\mathcal{O}}u(Y,t)dY$. By the triangle inequality and the Poincaré inequality in the spatial variables,
(3.6) $\iint_{J_{r}\cap\Omega}|u(X,t)-c_{u}|dXdt\leq\iint_{J_{r}\cap\Omega}|u(X,t)-\langle u\rangle(t)|dXdt+\iint_{J_{r}\cap\Omega}|\langle u\rangle(t)-c_{u}|dXdt\\ \leq Cr\int_{-r^{2}}^{0}\int_{Q^{\prime}\cap\mathcal{O}}\left|\nabla u(X,t)\right|dXdt+|Q^{\prime}\cap\mathcal{O}|\int_{-r^{2}}^{0}|\langle u\rangle(t)-c_{u}|dt.$
@@
$\langle u\rangle_{\eta}(t):=\int_{\mathcal{O}}u(Y,t)\eta(Y)dY\quad\text{for }t\in(T_{0},T_{1}),$
and set $c_{\eta}:=\int_{\mathcal{O}}\eta\,dX$. For $t,t^{\prime}\in[-r^{2},0]$, we have
@@
since $u$ is a weak solution to $Lu=-\operatorname{div}F$ with zero Neumann data on $\partial\Omega\cap Q_{0}$ (see (2.26) and the remark after it). Using the bound on $\nabla\eta$, we obtain that
(3.7) $|\langle u\rangle_{\eta}(t)-\langle u\rangle_{\eta}(t^{\prime})|\leq\frac{C}{r}\int_{\mathcal{O}\cap 2Q^{\prime}}\int_{-r^{2}}^{0}\left(|\nabla u|+|F|\right)dtdX\quad\text{for }t,t^{\prime}\in[-r^{2},0].$
@@
$\int_{-r^{2}}^{0}|\langle u\rangle(t)-c_{u}|dt=\int_{-r^{2}}^{0}\left|\fint_{-r^{2}}^{0}\left(\langle u\rangle(t)-\langle u\rangle(t^{\prime})\right)dt^{\prime}\right|dt\leq Cr\fint_{\mathcal{O}\cap 2Q^{\prime}}\int_{-r^{2}}^{0}\left(|\nabla u|+|F|\right)dtdX.$
@@
$\iint_{J_{r}\cap\Omega}|u(X,t)-c_{u}|dXdt\leq Cr\int_{\mathcal{O}\cap 2Q^{\prime}}\int_{-r^{2}}^{0}\left(|\nabla u|+|F|\right)dtdX,$
as desired. ∎
In the following lemma we show relations between the Dirichlet and Poisson-Dirichlet problems. We remark that although it can be shown that the solvability of $\rm{(}\rm{D}_{p^{\prime}}\rm{)}$$^{L^{*}}$, $\rm{(}\rm{PD}_{p^{\prime}}\rm{)}$$^{L^{*}}$ and $\rm{(}\rm{PR}_{p}\rm{)}$L are actually equivalent, we will only derive the partial implications that are sufficient for our purpose.
@@
by the assumption that $\rm{(}\rm{PD}_{p^{\prime}}\rm{)}$$^{L^{*}}$ is solvable. Again by (2.14), we have
@@
(4.3) $\|\widetilde{N}(\nabla u)\|_{L^{p}(\partial\Omega)}\leq\|\widetilde{N}(\nabla u_{D})\|_{L^{p}(\partial\Omega)}+\|\widetilde{N}(\nabla w)\|_{L^{p}(\partial\Omega)}\\ \leq\|\widetilde{N}(\nabla u_{D})\|_{L^{p}(\partial\Omega)}+\|g\|_{L^{p}(\partial\Omega)}\leq 2\|\widetilde{N}(\nabla u_{D})\|_{L^{p}(\partial\Omega)}\lesssim\|\widetilde{\mathcal{C}}(F)\|_{L^{p}(\partial\Omega)}.$
@@
where we have used that $\|F\|_{L^{2}(\Omega)}<\infty$ since $F\in L_{c}^{\infty}(\Omega)$. Hence, there exists a unique $g\in\dot{H}{\vphantom{H}}^{-1/4}_{\partial_{t}-\Delta_{x}}(\partial\Omega)$ such that $\Lambda(\varphi)=\langle g,\varphi\rangle$ for all $\varphi\in\dot{H}^{1/4}_{\partial_{t}-\Delta_{x}}(\partial\Omega)$. As Lipschitz functions $\varphi:\mathbb{R}^{n+1}\to\mathbb{R}$ with bounded support restricted to $\partial\Omega$ are dense in $\dot{H}^{1/4}_{\partial_{t}-\Delta_{x}}(\partial\Omega)$ it follows that the identity $\Lambda(\varphi)=\langle g,\varphi\rangle$ holds for all such $\varphi$ (and we can take $\Phi$ to be their Lipschitz extension in (4.5)) and still uniquely determined $g$ in the space $\dot{H}{\vphantom{H}}^{-1/4}_{\partial_{t}-\Delta_{x}}(\partial\Omega)$.
@@
To show the weak convergence of $g_{r}$ to $g$ in $L^{p}(\partial\Omega)$, it suffices to show that for any Lipschitz function $\varphi:\mathbb{R}^{n+1}\to\mathbb{R}$ with bounded support, we have that
@@
(4.14) $I\leq\iint_{\Omega}u\,Lv\,dX=\iint_{\Omega}u\left(-\partial_{t}v+\operatorname{div}A\nabla v\right)dXdt=\iint_{\Omega}\left(\partial_{t}u\,v-A^{T}\nabla u\cdot\nabla v\right)\,dXdt\\ =\iint_{\Omega}F\cdot\nabla v\,dXdt$
because $u$ is a weak solution to $L^{*}u=-\mathrm{div}F$ with zero Neumann data. By the Carleson inequality (2.14), the assumption that $\rm{(}\rm{wPNR}_{p}\rm{)}$ holds for $v$, and (4.13), we have that
@@
Let $p\in(1,\infty)$. Suppose that $\rm{(}\rm{wPN}_{p^{\prime}}\rm{)}$$^{L^{*}}$ is solvable in $\Omega=\mathcal{O}\times\mathbb{R}$. Let $Q_{0}:=Q_{0}^{\prime}\times(T_{0},T_{1})$ be a cube centered at $\partial\Omega$. Then for any $(\bar{x},\bar{t})\in Q_{0}\cap\partial\Omega$, any backward parabolic cube $J_{r}(\bar{x},\bar{t}):=Q_{r}^{\prime}(\bar{x})\times(\bar{t}-r^{2},\bar{t})$ that satisfies $J_{2r}(\bar{x},\bar{t})\subset Q_{0}$, and any weak solution $u$ to $Lu=0$ in $Q_{0}\cap\Omega$ satisfying $u\in L^{2}\left((T_{0},T_{1});W^{1,2}(Q_{0}^{\prime}\cap\mathcal{O})\right)$, $\partial_{t}u\in L^{2}\left((T_{0},T_{1});W^{-1,2}(Q_{0}^{\prime}\cap\mathcal{O})\right)$ and having zero Neumann data on $Q_{0}\cap\partial\Omega$, we have
(4.16) $\|\widetilde{N}(|\nabla u|{\mathds{1}}_{J_{r}(\bar{x},\bar{t})})\|_{L^{p}(\partial\Omega)}\leq Cr^{(n+1)/p}\left(\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{J_{2r}(\bar{x},\bar{t})\cap\Omega}|\nabla u|^{2}dXdt\right)^{\frac{1}{2}}.$